# Optimal control under unknown intensity with Bayesian learning

### v1

@@ 4.3 Comparison theorem
    $\quad+\alpha((n+1)-n).$
    $\leq\alpha.$
    We used crucially that $F$ is non-decreasing and that $|\bar{x}_{2}|=\lVert\bar{x}^{*}\rVert\leq\lVert\bar{x}\rVert$, as just after a jump, the potential is reset to zero.
@@
    For $\tilde{k}(\alpha,\epsilon,\delta):=k(\alpha,\epsilon,\delta)+2\lVert w\rVert_{\infty}L\sqrt{2C\epsilon}+\alpha$, we have $\beta\leq\tilde{k}(\alpha,\epsilon,\delta)$. As
@@
    there exists $\alpha(\beta)>0$, $\epsilon(\beta,\alpha)>0$ and $\delta(\beta,\alpha,\epsilon)$ such that
@@
    This choice is a contraction with $\beta\leq\tilde{k}(\alpha,\epsilon,\delta)$. Therefore, for those values we necessarily have $\bar{t}=T$ or $\bar{t^{\prime}}=T$. We conclude that
@@
    However, when $\epsilon$ is small enough, this contradicts (4.7). This ends the proof.
@@ 5 Examples
To illustrate numerically the results, we consider the drift $b(y)=-y$. The controlled process is:
@@
We present two examples using different prior distributions and intensity functions. We first consider an intensity function of the form
@@
where $\sum_{i=1}^{k}p_{i}=1$ and $k\geq 2$. In order to satisfy the assumptions, the function $f_{\lambda}$ can be artificially bounded by a constant $C>0$ so that it does not play an essential role. It is noteworthy that this family of priors is conjugate: the posterior distribution retains the same form, with updated weights $(p_{1},\ldots,p_{k})$. However, within our framework, the dimensionality of the problem is reduced to 2, as opposed to the original $k-1$.
We solve numerically the PDE, using a standard explicit scheme. The analysis of the convergence of this explicit scheme towards the unique viscosity solution of (4.1) can be done using the methodology of [3, Th. 2.1.].
Figure 1: An optimal trajectory with intensity $\lambda\exp(2(y-1))$ and with the prior distribution $\frac{1}{10}\left(\delta_{0}+2\delta_{0.25}+4\delta_{0.5}+2\delta_{0.75}+\delta_{1}\right)$. True value is $\lambda=1$.
In Figure 1, we present a simulated path of the optimal strategy, where the initial measure is given by $m_{0}:=\frac{1}{10}\left(\delta_{0}+2\delta_{0.25}+4\delta_{0.5}+2\delta_{0.75}+\delta_{1}\right)$ and the true parameter value is $\Lambda=1$. The top graphic illustrates the evolution of the potential over time. The second graphic depicts the optimal control, while the third and fourth graphics show the mean and variance of the posterior distribution, respectively. Finally, the bottom-left graphic displays $m_{0}$ as a histogram, and the bottom-right graphic presents $m_{T}$.
Throughout the trajectory, the controller applies a control $t\mapsto\gamma_{t}$ that stays close to one. Near the terminal date, after the final jump, the control becomes small, as increasing the potential further is not worth the cost.
Although the prior is centered at $\lambda=0.5$ with a low probability assigned to $\lambda=1$, the mean of the posterior distribution increases on average. By the end, the posterior distribution assigns the highest probabilities to $\lambda=0.75$ and $\lambda=1$, while excluding $\lambda=0$ and almost completely ruling out $\lambda=0.25$. It is important to note that the posterior probability of $\lambda=0$ immediately becomes zero as soon as a jump is observed.
For the second example, we modify the intensity function as follows:
@@
This function provides a continuous approximation to the discontinuous function $(\lambda,y)\mapsto\lambda\mathbf{1}_{\{x\geq 1\}}$. In the case of the discontinuous function, for $x<1$, the controller must decide whether to increase the potential to 1 or not. Once the potential reaches 1, it remains fixed thereafter. The continuous version introduced here should exhibit similar overall behavior. For the prior distribution, we choose the uniform probability measure on $[0,2]$:
@@
Figure 2: An optimal trajectory with intensity close to $\lambda\mathbf{1}_{\{y\geq 1\}}$ and the prior $\mathcal{U}\left([0,2]\right)$. True value is $\lambda=1$.
In Figure 2, we present a simulated path of the optimal strategy, where the true parameter value is $\Lambda=1$. The graphics illustrate the same functions as in Figure 1, with the distinction that the bottom-left and bottom-right panels display $m_{0}$ and $m_{T}$ as density functions, respectively.
Throughout the trajectory, the controller applies a control $t\mapsto\gamma_{t}$ that increases until the potential reaches a value slightly larger than $Y_{t}=1$. At this point, the control makes a jump to a value close to $\gamma_{t}=1$. Since the drift of the potential is given by $-Y_{t}+\gamma_{t}$, setting $\gamma_{t}=1$ once the potential reaches 1 is optimal, as there is no incentive to exceed this value. As the terminal date approaches, following the final jump, the control value decreases; further increases in potential become unfeasible due to the associated costs.
The prior begins centered on the true value. While the potential remains below one, it remains almost unchanged since we cannot obtain information about $\Lambda$. Once it reaches one, both the mean and variance decrease linearly. When a jump is observed, the mean increases, while the behavior of the variance depends on the specifics of the observation. However, the variance generally decreases over time. Although the prior is symmetric at $\lambda=1$, the observation of a jump excludes the value $\lambda=0$, leading to a density of zero at that point. In contrast, the density at the upper boundary, $\lambda=2$, remains positive, nonetheless close to zero.
@@ Acknowledgments
Nicolas Baradel acknowledges the financial support provided by the Fondation Natixis. Quentin Cormier thanks Carl Graham for useful discussions at different stage of this project, as well as Charles Bertucci and Mete Soner for their help with viscosity solutions.

### v2

@@ 4.3 Comparison theorem
    $\quad+\alpha((n+1)-n)$
    $\leq\alpha,$
    using that $F$ is non-decreasing and that $|\bar{x}_{2}|=\lVert\bar{x}^{*}\rVert\leq\lVert\bar{x}\rVert$, since the $y$-coordinate is reset to zero after a jump.
@@
    For $\tilde{k}(\alpha,\epsilon,\delta):=k(\alpha,\epsilon,\delta)+2\lVert w\rVert_{\infty}L\sqrt{2C\epsilon}+\alpha$, we have $\beta\leq\tilde{k}(\alpha,\epsilon,\delta)$. Since:
@@
    there exists $\alpha(\beta)>0$, $\epsilon(\beta,\alpha)>0$ and $\delta(\beta,\alpha,\epsilon)$ such that:
@@
    This choice is in contraction with $\beta\leq\tilde{k}(\alpha,\epsilon,\delta)$. Thus, either $\bar{t}=T$ or $\bar{t^{\prime}}=T$, and:
@@
    For $\epsilon$ sufficiently small, this contradicts (4.7), completing the proof.
@@ 5 Examples
To illustrate the results numerically, we consider the drift $b(y)=-y$. The controlled process is:
@@
We present two examples using different prior distributions and intensity functions.
5.1 First example
We first consider an intensity function of the form:
@@ 5.1 First example
where $\sum_{i=1}^{k}p_{i}=1$ and $k\geq 2$. To satisfy Assumption 3.1, the function $f_{\lambda}$ can be artificially bounded by a constant $C>0$, which does not significantly affect the results. Notably, this family of priors is conjugate: the posterior distribution retains the same form with updated weights $(p_{1},\ldots,p_{k})$. Within our framework, the dimensionality of the problem is reduced to two, as opposed to the original $k-1$.
We solve the PDE (4.1) numerically using a standard explicit scheme. The convergence of this scheme to the unique viscosity solution can be analyzed following [3, Thm. 2.1].
Figure 1: An optimal trajectory with intensity $\lambda\exp(2(y-1))$ and prior distribution $\frac{1}{10}\left(\delta_{0}+2\delta_{0.25}+4\delta_{0.5}+2\delta_{0.75}+\delta_{1}\right)$. The true value is $\lambda=1$.
In Figure 1, we present a simulated path of the optimal strategy, where the initial measure is $m_{0}:=\frac{1}{10}\left(\delta_{0}+2\delta_{0.25}+4\delta_{0.5}+2\delta_{0.75}+\delta_{1}\right)$ and the true parameter value is $\Lambda=1$. The top plot shows the evolution of the potential over time. The second plot depicts the optimal control, while the third and fourth plots display the mean and variance of the posterior distribution, respectively. The bottom-left plot shows $m_{0}$ as a histogram, and the bottom-right plot presents $m_{T}$.
Throughout the trajectory, the controller applies a control $t\mapsto\gamma_{t}$ that remains close to 1. Near the terminal time, after the final jump, the control decreases, as further increasing the potential is not cost-effective.
Although the prior is centered at $\lambda=0.5$ with low probability assigned to $\lambda=1$, the mean of the posterior distribution increases on average. By the end, the posterior assigns the highest probabilities to $\lambda=0.75$ and $\lambda=1$, excluding $\lambda=0$ and nearly ruling out $\lambda=0.25$. Notably, the posterior probability of $\lambda=0$ becomes zero immediately after observing a jump.
5.2 Second example
For the second example, we consider the intensity function:
@@ 5.2 Second example
This function approximates the discontinuous function $(\lambda,y)\mapsto\lambda\mathbf{1}_{\{y\geq 1\}}$. For the discontinuous case, when $y<1$, the controller decides whether to increase the potential to 1. Once the potential reaches 1, it remains fixed. The continuous version should exhibit similar behavior. We choose the uniform prior distribution on $[0,2]$:
@@
Figure 2: An optimal trajectory with intensity approximating $\lambda\mathbf{1}_{\{y\geq 1\}}$ and prior $\mathcal{U}\left([0,2]\right)$. The true value is $\lambda=1$.
In Figure 2, we present a simulated path of the optimal strategy with $\Lambda=1$. The plots illustrate the same quantities as in Figure 1, except that the bottom-left and bottom-right panels display $m_{0}$ and $m_{T}$ as density functions, respectively.
Throughout the trajectory, the controller applies a control $t\mapsto\gamma_{t}$ that increases until the potential reaches slightly above $Y_{t}=1$ at which point the control jumps to approximately $\gamma_{t}=1$. Since the drift of the potential is $-Y_{t}+\gamma_{t}$, setting $\gamma_{t}=1$ when the potential reaches 1 is optimal, as exceeding this value is unnecessary. Near the terminal time, after the final jump, the control decreases, as further increases in potential are not cost-effective.
The prior is initially centered at the true value. While the potential remains below 1, the posterior distribution is nearly unchanged, as no information about $\Lambda$ is gained. Once the potential reaches 1, the mean and variance decrease linearly. Upon observing a jump, the mean increases, while the variance’s behavior depends on the observation specifics but generally decreases over time. Although the prior is symmetric around $\lambda=1$, a jump excludes $\lambda=0$, resulting in zero density at that point. The density at the upper boundary, $\lambda=2$, remains positive but close to zero.
@@ Acknowledgments
Nicolas Baradel acknowledges the financial support from the Fondation Natixis. Quentin Cormier thanks Carl Graham for insightful discussions at various stage of this project, as well as Charles Bertucci and Mete Soner for their help with viscosity solutions.